Show that the function f defined by


f(x,y) = {1, (x,y) = (0,0)

{(x2 + y)/(x + y), (x,y) ≠ (0,0)


is not continuous by using two path test at (0,0).


1
Expert's answer
2022-05-12T03:20:53-0400

If "x=0" then "f(0, y)=\dfrac{0+y}{0+y}=1." Therefore


"f(x, y)\to 1\text{ as} \ (x,y)\to(0,0) \text{ along the }y-\text{axis}"

For all "x\not=0"


"f(x, x)=\dfrac{x^2+x}{x+x}=\dfrac{x+1}{2}"

"f(x, y)\to \dfrac{1}{2}\not=1\text{ as} \ (x,y)\to(0,0) \text{ along }y=x"

Since we have obtained different limits along different paths, limit


"\lim\limits_{(x,y)\to(0,0)}f(x, y)=\lim\limits_{(x,y)\to(0,0)}\dfrac{x^2+y}{x+y}"

does not exist.

The function

"f(x, y)= \begin{cases}
 1, & (x,y)=(0,0) \\
 \dfrac{x^2+y}{x+y}, &(x,y)\not=(0,0)
\end{cases}"

is not continuous at "(0,0)."